# Avoiding Catastrophe in Online Learning by Asking for Help

### v4

@@ A.2 Proof
Lemma 5.4 is a restatement and simplification of Lemma 3.5 in Russo et al. (2024). First, Russo et al. (2024) parametrize their algorithm by the expected number of queries $\hat{k}$ instead of the query probability $p=\hat{k}/T$. Second, Russo et al. (2024) include a second parameter $k$, which is the eventual target number of queries for their unconditional query bound. In our case, an expected query bound is sufficient, so we simply set $k=\hat{k}$. Third, Russo et al. (2024) provide a second bound which is tighter for small $k$; that bound is less useful for us so we omit it. Fourth, their number of actions $n$ is equal to $|\tilde{\Pi}|$ in our setting. (Their actions correspond to policies in $\tilde{\Pi}$, not our actions in $\mathcal{Y}$.) Since Russo et al. (2024) set $\eta=\max\Big{(}\frac{1}{T}\sqrt{\frac{\hat{k}\log n}{2}},\frac{k\hat{k}}{\sqrt{2}T^{2}}\Big{)}$, we end up with $\eta=\max\Big{(}\sqrt{\frac{p\log|\tilde{\Pi}|}{2T}},\>\frac{p^{2}}{\sqrt{2}}\Big{)}$. Algorithm 2 provides precise pseudocode for the HedgeWithQueries algorithm to which Lemma 5.4 refers.
1:function HedgeWithQueries($p\in(0,1],\>\text{finite policy class }\tilde{\Pi},\>\text{unknown }\ell:[T]\times\tilde{\Pi}\to[0,1]$)
2:     $w(\pi)\leftarrow 1$ for all $\pi\in\tilde{\Pi}$
3:     $\eta\leftarrow\max\big{(}\sqrt{\frac{p\log|\tilde{\Pi}|}{2T}},\>\frac{p^{2}}{\sqrt{2}}\big{)}$
4:     for $t$ from $1$ to $T$ do
@@
6:         if hedgeQuery then
7:              Query and observe $\ell(t,\pi)$ for all $\pi\in\tilde{\Pi}$
8:              $\ell^{*}\leftarrow\min_{\pi\in\tilde{\Pi}}\ell(t,\pi)$
9:              $w(\pi)\leftarrow w(\pi)\cdot\exp(-\eta(\ell(t,\pi)-\ell^{*}))$ for all $\pi\in\tilde{\Pi}$
10:              Select policy $\arg\min_{\pi\in\tilde{\Pi}}\ell(t,\pi)$
11:         else
12:              $P(\pi)\leftarrow w(\pi)/\sum_{\pi^{\prime}\in\tilde{\Pi}}w(\pi^{\prime})$ for all $\pi\in\tilde{\Pi}$
13:              Sample $\pi_{t}\sim P$
14:              Select policy $\pi_{t}$
@@ B.2 Main proof
    For each $t\in[T]$, let $X_{t}$ refer to the value of $X$ at the start of time step $t$.
@@
    For $X\subseteq\mathcal{X}$, let $\operatorname*{vol}(X)$ denote the $n$-dimensional Lebesgue measure of $X$.
@@
    We will use “Case 1” to refer to finite VC dimension and $\sigma$-smooth $\boldsymbol{x}$ and “Case 2” to refer to finite Littlestone dimension. In Case 1, expectations are over the randomness of both $\boldsymbol{x}$ and the algorithm, while in Case 2, expectations are over just the randomness of the algorithm. When we need to distinguish, we use $\operatorname*{\mathbb{E}}_{\boldsymbol{y}}$ to denote the expectation over randomness of the algorithm and $\operatorname*{\mathbb{E}}_{\boldsymbol{x}\sim\boldsymbol{\mathcal{D}}}$ to denote the expectation over $\boldsymbol{x}$.
@@
Under the conditions of Theorem 5.1, Algorithm 1 satisfies
$\operatorname*{\mathbb{E}}[|M_{T}|]\in O\left(\frac{d}{\sigma}T\varepsilon\log(1/\varepsilon)\log T\right)$
@@
Define $\ell:[T]\times\tilde{\Pi}\to[0,1]$ by $\ell(t,\pi)=\mathbf{1}(\pi(x_{t})\neq\pi^{m}(x_{t}))$, and let $w^{h}$ and $\pi_{t}^{h}$ denote the values of $w$ and $\pi_{t}$ respectively in HedgeWithQueries, while $w$ and $\pi_{t}$ refer to the variables in Algorithm 1. Then $w$ and $w^{h}$ evolve in the exact same way, so the distributions of $\pi_{t}$ and $\pi_{t}^{h}$ coincide. Thus by Lemma 5.4,
$\operatorname*{\mathbb{E}}_{\boldsymbol{y}}\left[\sum_{t=1}^{T}\ell(t,\pi_{t})\right]-\min_{\tilde{\pi}\in\tilde{\Pi}}\sum_{t=1}^{T}\ell(t,\tilde{\pi})\leq\frac{2\log|\tilde{\Pi}|}{p^{2}}=2T\varepsilon\log|\tilde{\Pi}|$
Since Lemma 5.4 holds for any loss function, the bound above holds for any $\boldsymbol{x}\in S^{T}$, so the bound also holds in expectation over $\boldsymbol{x}\sim\boldsymbol{\mathcal{D}}$ (which is needed for Case 1). Next, observe that $|M_{T}|=\sum_{t=1}^{T}\mathbf{1}(\pi_{t}(x_{t})\neq\pi^{m}(x_{t}))=\sum_{t=1}^{T}\ell(t,\pi_{t})$, so
$\operatorname*{\mathbb{E}}_{\boldsymbol{y}}[|M_{T}|]\leq 2T\varepsilon\log|\tilde{\Pi}|+\min_{\tilde{\pi}\in\tilde{\Pi}}\sum_{t=1}^{T}\mathbf{1}(\tilde{\pi}(x_{t})\neq\pi^{m}(x_{t}))$
Case 1: Since $\tilde{\Pi}$ is a smooth $\varepsilon$-cover of $\Pi$, we have
$\operatorname*{\mathbb{E}}_{\boldsymbol{x}\sim\boldsymbol{\mathcal{D}}}\left[\min_{\tilde{\pi}\in\tilde{\Pi}}\sum_{t=1}^{T}\mathbf{1}(\tilde{\pi}(x_{t})\neq\pi^{m}(x_{t}))\right]\leq$ $\ \operatorname*{\mathbb{E}}_{\boldsymbol{x}\sim\boldsymbol{\mathcal{D}}}\left[\sup_{\pi\in\Pi}\min_{\tilde{\pi}\in\tilde{\Pi}}\sum_{t=1}^{T}\mathbf{1}(\tilde{\pi}(x_{t})\neq\pi(x_{t}))\right]$
$\in$ $\ O\left(\frac{1}{\sigma}T\varepsilon\log T\sqrt{d\log(1/\varepsilon)}\right)$
with the first step due to $\pi^{m}\in\Pi$ and the second step due to Lemma 5.3. The last component we need is that $|\tilde{\Pi}|\leq(41/\varepsilon)^{d}$ by construction (and such a $\tilde{\Pi}$ is guaranteed to exist by Lemma 5.1). Combining the above inequalities and taking the expectation over $\boldsymbol{x}\sim\boldsymbol{\mathcal{D}}$ (in addition to the randomness of the algorithm), we get
$\operatorname*{\mathbb{E}}_{\boldsymbol{x}\sim\boldsymbol{\mathcal{D}},\,\boldsymbol{y}}\,[|M_{T}|]\leq$ $\ 2T\varepsilon\log|\tilde{\Pi}|+\operatorname*{\mathbb{E}}_{\boldsymbol{x}\sim\boldsymbol{\mathcal{D}}}\left[\min_{\tilde{\pi}\in\tilde{\Pi}}\sum_{t=1}^{T}\mathbf{1}(\tilde{\pi}(x_{t})\neq\pi^{m}(x_{t}))\right]$
$\leq$ $\ 2dT\varepsilon\log(41/\varepsilon)+O\left(\frac{1}{\sigma}T\varepsilon\log T\sqrt{d\log(1/\varepsilon)}\right)$
$\in$ $\ O\left(\frac{d}{\sigma}T\varepsilon\log(1/\varepsilon)\log T\right)$
@@
$\operatorname*{\mathbb{E}}_{\boldsymbol{y}}[|M_{T}|]\leq$ $\ 2T\varepsilon\log|\tilde{\Pi}|+\min_{\tilde{\pi}\in\tilde{\Pi}}\sum_{t=1}^{T}\mathbf{1}(\tilde{\pi}(x_{t})\neq\pi^{m}(x_{t}))$
$\leq$ $\ 2T\varepsilon d\ln(eT/d)$
$\in$ $\ O\left(\frac{d}{\sigma}T\varepsilon\log(1/\varepsilon)\log T\right)$
@@
Consider an arbitrary $t\in[T]$. If $t\in Q_{T}$, then $\mu_{t}(x_{t},y_{t})=\mu_{t}^{m}(x_{t})$, so assume $t\not\in Q_{T}$. Let $(x^{\prime},y^{\prime})=\operatorname*{arg\,min}_{(x,y)\in X_{t}:\pi_{t}(x_{t})=y}||x_{t}-x||$. Since $t\not\in Q_{T}$, we must have $||x_{t}-x^{\prime}||\leq\varepsilon^{1/n}$.
We have $y^{\prime}=\pi^{m}(x^{\prime})$ by construction of $X_{t}$ and $\pi_{t}(x_{t})=y^{\prime}$ by construction of $y^{\prime}$. Combining these with the local generalization assumption, we get
$\mu_{t}(x_{t},y_{t})=\mu_{t}(x_{t},\pi_{t}(x_{t}))=\mu_{t}(x_{t},\pi^{m}(x^{\prime}))\geq\mu_{t}^{m}(x_{t})-L||x_{t}-x^{\prime}||\geq\mu_{t}^{m}(x_{t})-L\varepsilon^{1/n}$
@@
Assume $a_{1},\dots,a_{T},b_{1},\dots,b_{T}\in[0,1]$ and $a_{t}\geq b_{t}$ for all $t\in[T]$. Then
$\prod_{t=1}^{T}a_{t}-\prod_{t=1}^{T}b_{t}\leq\sum_{t=1}^{T}a_{t}-\sum_{t=1}^{T}b_{t}$
@@
We proceed by induction on $T$. The claim is trivially satisfied for $T=1$, so suppose $T>1$ and assume that $\prod_{t=1}^{T-1}a_{t}-\prod_{t=1}^{T-1}b_{t}\leq\sum_{t=1}^{T-1}a_{t}-\sum_{t=1}^{T-1}b_{t}$. Then
$\sum_{t=1}^{T}a_{t}-\sum_{t=1}^{T}b_{t}-\prod_{t=1}^{T}a_{t}+\prod_{t=1}^{T}b_{t}=$ $\ a_{T}\sum_{t=1}^{T-1}a_{t}-b_{T}\sum_{t=1}^{T-1}b_{t}-a_{T}\prod_{t=1}^{T-1}a_{t}+b_{T}\prod_{t=1}^{T-1}b_{t}$
$=$ $\ a_{T}\left(\sum_{t=1}^{T-1}a_{t}-\prod_{t=1}^{T-1}a_{t}\right)-b_{T}\left(\sum_{t=1}^{T-1}b_{t}-\prod_{t=1}^{T-1}b_{t}\right)$
Since $T>1$ and $a_{t}\in[0,1]$ for all $t\in[T]$, we have $\sum_{t=1}^{T-1}a_{t}\geq a_{1}\geq\sum_{t=1}^{T-1}a_{t}$. Thus $\sum_{t=1}^{T-1}a_{t}-\prod_{t=1}^{T-1}a_{t}\geq 0$. Combining this with $a_{T}\geq b_{T}$, we get
$\sum_{t=1}^{T}a_{t}-\sum_{t=1}^{T}b_{t}-\prod_{t=1}^{T}a_{t}+\prod_{t=1}^{T}b_{t}=$ $\ a_{T}\left(\sum_{t=1}^{T-1}a_{t}-\prod_{t=1}^{T-1}a_{t}\right)-b_{T}\left(\sum_{t=1}^{T-1}b_{t}-\prod_{t=1}^{T-1}b_{t}\right)$
$\geq$ $\ b_{T}\left(\sum_{t=1}^{T-1}a_{t}-\prod_{t=1}^{T-1}a_{t}\right)-b_{T}\left(\sum_{t=1}^{T-1}b_{t}-\prod_{t=1}^{T-1}b_{t}\right)$
$=$ $\ b_{T}\left(\sum_{t=1}^{T-1}a_{t}-\prod_{t=1}^{T-1}a_{t}-\sum_{t=1}^{T-1}b_{t}+\prod_{t=1}^{T-1}b_{t}\right)$
@@
The last step is due to $b_{T}\geq 0$ and our assumption of $\prod_{t=1}^{T-1}a_{t}-\prod_{t=1}^{T-1}b_{t}\leq\sum_{t=1}^{T-1}a_{t}-\sum_{t=1}^{T-1}b_{t}$. ∎
Lemma B.4.
Under the conditions of Theorem 5.1, Algorithm 1 satisfies
$\operatorname*{\mathbb{E}}\left[R_{T}\right]\in$ $\ O\left(\frac{dL}{\sigma}T\varepsilon^{1+1/n}\log(1/\varepsilon)\log T\right)$
$\operatorname*{\mathbb{E}}\left[\sum_{t=1}^{T}\mu_{t}^{m}(x_{t})-\sum_{t=1}^{T}\mu_{t}(x_{t},y_{t})\right]\in$ $\ O\left(\frac{dL}{\sigma}T\varepsilon^{1+1/n}\log(1/\varepsilon)\log T\right)$
Proof.
We first claim that $y_{t}=\pi^{m}(x_{t})$ for all $t\not\in M_{T}$. If $t\in Q_{T}$, the claim is immediate; if not, we have $y_{t}=\pi_{t}(x_{t})$, and $\pi_{t}(x_{t})=\pi^{m}(x_{t})$ due to $t\not\in M_{T}$. Thus $\min(\mu_{t}^{m}(x_{t}),\mu_{t}(x_{t},y_{t}))=\mu_{t}^{m}(x_{t})$ for $t\not\in M_{T}$.
We next claim that $\mu_{t}^{m}(x_{t})-\min(\mu_{t}^{m}(x_{t}),\mu_{t}(x_{t},y_{t}))\leq L\varepsilon^{1/n}$ for all $t\in[T]$. If $\mu_{t}(x_{t},y_{t})\leq\mu_{t}^{m}(x_{t})$, this follows from Lemma B.2. If $\mu_{t}(x_{t},y_{t})>\mu_{t}^{m}(x_{t})$, then $\mu_{t}^{m}(x_{t})-\min(\mu_{t}^{m}(x_{t}),\mu_{t}(x_{t},y_{t}))=0\leq L\varepsilon^{1/n}$. Therefore
$\sum_{t=1}^{T}\Big{(}\mu_{t}^{m}(x_{t})-\min(\mu_{t}^{m}(x_{t}),\mu_{t}(x_{t},y_{t}))\Big{)}\leq$ $\ \sum_{t\in M_{T}}\Big{(}\mu_{t}^{m}(x_{t})-\min(\mu_{t}^{m}(x_{t}),\mu_{t}(x_{t},y_{t}))\Big{)}$
$\leq$ $\ \sum_{t\in M_{T}}L\varepsilon^{1/n}$
$=$ $\ |M_{T}|L\varepsilon^{1/n}$
Now let $a_{t}=\mu_{t}^{m}(x_{t})$ and $b_{t}=\min(\mu_{t}^{m}(x_{t}),\mu_{t}(x_{t},y_{t}))$ for all $t\in[T]$. Then by Lemma B.3,
$\prod_{t=1}^{T}\mu_{t}^{m}(x_{t})-\prod_{t=1}^{T}\min(\mu_{t}^{m}(x_{t}),\mu_{t}(x_{t},y_{t}))\leq$ $\ \sum_{t=1}^{T}\Big{(}\mu_{t}^{m}(x_{t})-\min(\mu_{t}^{m}(x_{t}),\mu_{t}(x_{t},y_{t}))\Big{)}$
Since $\mu_{t}(x_{t},y_{t})\geq\min(\mu_{t}^{m}(x_{t}),\mu_{t}(x_{t},y_{t}))$ for all $t\in[T]$, we have
$R_{T}=$ $\ \prod_{t=1}^{T}\mu_{t}^{m}(x_{t})-\prod_{t=1}^{T}\mu_{t}(x_{t},y_{t})$
$\leq$ $\ \sum_{t=1}^{T}\Big{(}\mu_{t}^{m}(x_{t})-\min(\mu_{t}^{m}(x_{t}),\mu_{t}(x_{t},y_{t}))\Big{)}$
$\leq$ $\ |M_{T}|L\varepsilon^{1/n}$
Since we also have $\sum_{t=1}^{T}\mu_{t}^{m}(x_{t})-\sum_{t=1}^{T}\mu_{t}(x_{t},y_{t})\leq\sum_{t=1}^{T}(\mu_{t}^{m}(x_{t})-\min(\mu_{t}^{m}(x_{t}),\mu_{t}(x_{t},y_{t})))$, we get both of the following
$R_{T}\leq$ $\ L\varepsilon^{1/n}|M_{T}|$
$\sum_{t=1}^{T}\mu_{t}^{m}(x_{t})-\sum_{t=1}^{T}\mu_{t}(x_{t},y_{t})\leq$ $\ L\varepsilon^{1/n}|M_{T}|$
Taking the expectation and applying Lemma B.1 to both of the above inequalities completes the proof. ∎
@@
Let $(K,||\cdot||)$ be a normed vector space and let $\delta>0$. Then $X\subseteq K$ is a $\delta$-packing of $K$ if for all $x,y\in X$, $||x-y||>\delta$. The $\delta$-packing number of $K$, denoted $\mathcal{M}(K,||\cdot||,\delta)$, is the maximum cardinality of any $\delta$-packing of $K$.
In this paper, we only consider the Euclidean distance norm, so we just write $M(K,||\cdot||,\delta)=M(K,\delta)$.
Lemma B.5 (Theorem 14.2 in Wu (2020)).
@@
[…]
@@
$\operatorname*{\mathbb{E}}[|Q_{T}|]\in O\left(\sqrt{\frac{T}{\varepsilon}}+\frac{d}{\sigma}T\varepsilon\log(1/\varepsilon)\log T+\frac{\operatorname*{diam}(\boldsymbol{x})^{n}}{\varepsilon}\right)$
@@
Next, fix an $y\in\mathcal{Y}$ and let $X_{y}=\{x\in\boldsymbol{x}:\pi^{m}(x)=y\}$ be the set of observed inputs which share a mentor action. We claim that $\hat{X}_{2}\cap X_{y}$ is a packing of $X_{y}$. Suppose instead that there exists $x,x^{\prime}\in\hat{X}_{2}\cap X_{y}$, with $||x-x^{\prime}||\leq\varepsilon^{1/n}$. WLOG assume $x$ was queried after $x^{\prime}$ and let $t$ be the time step on which $x$ was queried. Then $(x^{\prime},\pi^{m}(x^{\prime}))\in X_{t}$. Also, $x,x^{\prime}\in\hat{X}_{2}\cap X_{y}$ implies that and $\pi_{t}(x_{t})=\pi^{m}(x_{t})=y=\pi^{m}(x^{\prime})$. Therefore
$\min_{(x^{\prime\prime},y^{\prime\prime})\in X_{t}:y^{\prime\prime}=\pi_{t}(x_{t})}||x_{t}-x^{\prime\prime}||\leq||x_{t}-x^{\prime}||\leq\varepsilon^{1/n}$
which contradicts $x_{t}\in\hat{X}$. Thus $\hat{X}_{2}\cap X_{y}$ is a $\varepsilon^{1/n}$-packing of $X_{y}$.
By Lemma B.6, there exists a ball $B_{1}$ of diameter $\operatorname*{diam}(\boldsymbol{x})\sqrt{\frac{n}{2(n+1)}}$ which contains $\boldsymbol{x}$. Let $R=\operatorname*{diam}(\boldsymbol{x})\sqrt{\frac{n}{8(n+1)}}$ denote the radius of $B_{1}$. Let $B_{2}$ be the ball with the same center as $B_{1}$ but with radius $\max(R,\varepsilon^{1/n})$. Since $X_{y}\subset\boldsymbol{x}\subset B_{1}\subset B_{2}$, $\hat{X}_{2}\cap X_{y}$ is also a $\varepsilon^{1/n}$-packing of $B_{2}$. Also, $B_{2}$ must contain a ball of radius $\varepsilon^{1/n}$, so Lemma B.5 implies that
@@
$=$ $\ \max\left(\operatorname*{diam}(\boldsymbol{x})^{n}\left(\frac{n}{8(n+1)}\right)^{n/2},\>\varepsilon\right)\frac{3^{n}}{\varepsilon}$

### v5

@@ A.3 Proof
Lemma 5.4 is a restatement and simplification of Lemma 3.5 in Russo et al. (2024) with the following differences:
  1. 1.
    They parametrize their algorithm by the expected number of queries $\hat{k}$ instead of the query probability $p=\hat{k}/T$.
  2. 2.
    They include a second parameter $k$, which is the eventual target number of queries for their unconditional query bound. In our case, an expected query bound is sufficient, so we simply set $k=\hat{k}$.
  3. 3.
    They provide a second bound which is tighter for small $k$; that bound is less useful for us so we omit it.
  4. 4.
    Their “actions” correspond to policies in our setting, not actions in $\mathcal{Y}$. Their number of actions $n$ corresponds to $|\tilde{\Pi}|$.
  5. 5.
    We include an expectation over both loss terms, while they only include an expectation over the agent’s loss. This is because an adaptive adversary may choose the loss function for time $t$ in a randomized manner. Since we eventually set $\ell(t,\pi)=\mathbf{1}(\pi(x_{t})\neq\pi^{m}(x_{t}))$, the randomization in $\ell$ corresponds to the randomization in $x_{t}$.
Altogether, since Russo et al. (2024) set $\eta=\max\Big{(}\frac{1}{T}\sqrt{\frac{\hat{k}\log n}{2}},\frac{k\hat{k}}{\sqrt{2}T^{2}}\Big{)}$, we end up with $\eta=\max\Big{(}\sqrt{\frac{p\log|\tilde{\Pi}|}{2T}},\>\frac{p^{2}}{\sqrt{2}}\Big{)}$. Algorithm 2 provides precise pseudocode for the HedgeWithQueries algorithm to which Lemma 5.4 refers.
  function HedgeWithQueries$(p\in(0,1],\>\text{finite policy class }\tilde{\Pi},\>\text{unknown }\ell:[T]\times\tilde{\Pi}\to[0,1])$
     $w(\pi)\leftarrow 1$ for all $\pi\in\tilde{\Pi}$
     $\eta\leftarrow\max\big{(}\sqrt{\frac{p\log|\tilde{\Pi}|}{2T}},\>\frac{p^{2}}{\sqrt{2}}\big{)}$
     for $t$ from $1$ to $T$ do
@@
[…]
        if hedgeQuery then
           Query and observe $\ell(t,\pi)$ for all $\pi\in\tilde{\Pi}$
           $\ell^{*}\leftarrow\min_{\pi\in\tilde{\Pi}}\ell(t,\pi)$
           $w(\pi)\leftarrow w(\pi)\cdot\exp(-\eta(\ell(t,\pi)-\ell^{*}))$ for all $\pi\in\tilde{\Pi}$
           Select policy $\arg\min_{\pi\in\tilde{\Pi}}\ell(t,\pi)$
        else
           $P(\pi)\leftarrow w(\pi)/\sum_{\pi^{\prime}\in\tilde{\Pi}}w(\pi^{\prime})$ for all $\pi\in\tilde{\Pi}$
           Sample $\pi_{t}\sim P$
           Select policy $\pi_{t}$
@@ B.2 Main proof
    For each $t\in[T]$, let $S_{t}$ refer to the value of $S$ at the start of time step $t$.
@@
    For $S\subseteq\mathcal{X}$, let $\operatorname*{vol}(S)$ denote the $n$-dimensional Lebesgue measure of $S$.
@@
    We will use “Case 1” to refer to finite VC dimension and $\sigma$-smooth $\boldsymbol{x}$ and “Case 2” to refer to finite Littlestone dimension.
@@
Let $\mathcal{Y}=\{0,1\}$. Assume $\pi^{m}\in\Pi$ where either (1) $\Pi$ has finite VC dimension $d$ and $\boldsymbol{x}$ is $\sigma$-smooth, or (2) $\Pi$ has finite Littlestone dimension $d$. Then for any $T\in\mathbb{N}$ and $\varepsilon\geq 1/T$, [^12] Note that this lemma omits the assumption of $\varepsilon\leq(\frac{\mu_{0}^{m}}{2L})^{n}$, since we do not need it for this lemma, and we would like to apply this lemma in the multi-action case without that assumption. Algorithm 1 satisfies
$\operatorname*{\mathbb{E}}[|M_{T}|]\in O\left(\frac{d}{\sigma}T\varepsilon\log(T+1/\varepsilon)\right)$
@@
Define $\ell:[T]\times\tilde{\Pi}\to[0,1]$ by $\ell(t,\pi)=\mathbf{1}(\pi(x_{t})\neq\pi^{m}(x_{t}))$, and let $w^{h}$ and $\pi_{t}^{h}$ denote the values of $w$ and $\pi_{t}$ respectively in HedgeWithQueries, while $w$ and $\pi_{t}$ refer to the variables in Algorithm 1. Then $w$ and $w^{h}$ evolve in the exact same way, so the distributions of $\pi_{t}$ and $\pi_{t}^{h}$ coincide. Also, $\varepsilon\geq 1/T$ implies that $p=1/\sqrt{\varepsilon T}\in(0,1]$. Thus by Lemma 5.4,
$\sum_{t=1}^{T}\operatorname*{\mathbb{E}}[\ell(t,\pi_{t})]-\min_{\tilde{\pi}\in\tilde{\Pi}}\sum_{t=1}^{T}\operatorname*{\mathbb{E}}[\ell(t,\tilde{\pi})]\leq$ $\ \frac{\log|\tilde{\Pi}|}{p^{2}}$
$=$ $\ T\varepsilon\log|\tilde{\Pi}|$
Since $|M_{T}|=\sum_{t=1}^{T}\mathbf{1}(\pi_{t}(x_{t})\neq\pi^{m}(x_{t}))=\sum_{t=1}^{T}\ell(t,\pi_{t})$, we have
$\operatorname*{\mathbb{E}}[|M_{T}|]\leq T\varepsilon\log|\tilde{\Pi}|+\min_{\tilde{\pi}\in\tilde{\Pi}}\sum_{t=1}^{T}\operatorname*{\mathbb{E}}[\mathbf{1}(\tilde{\pi}(x_{t})\neq\pi^{m}(x_{t}))]$
Case 1: Since $\tilde{\Pi}$ is a smooth $\varepsilon$-cover and $\pi^{m}\in\Pi$, Lemma 5.1 implies that $\operatorname*{\mathbb{E}}[\mathbf{1}(\tilde{\pi}(x_{t})\neq\pi^{m}(x_{t}))]\leq\varepsilon/\sigma$ for any $\tilde{\pi}\in\tilde{\Pi}$. Since $|\tilde{\Pi}|\leq(41/\varepsilon)^{d}$ by construction (and such a $\tilde{\Pi}$ is guaranteed to exist by Lemma 5.2), we get
$\operatorname*{\mathbb{E}}[|M_{T}|]\leq$ $\ T\varepsilon\log((41/\varepsilon)^{d})+\min_{\tilde{\pi}\in\tilde{\Pi}}\sum_{t=1}^{T}\frac{\varepsilon}{\sigma}$
$=$ $\ dT\varepsilon\log(41/\varepsilon)+\frac{T\varepsilon}{\sigma}$
$\in$ $\ O\left(\frac{d}{\sigma}T\varepsilon\log(T+1/\varepsilon)\right)$
@@
$\operatorname*{\mathbb{E}}[|M_{T}|]\leq$ $\ T\varepsilon\log|\tilde{\Pi}|+\min_{\tilde{\pi}\in\tilde{\Pi}}\sum_{t=1}^{T}\mathbf{1}(\tilde{\pi}(x_{t})\neq\pi^{m}(x_{t}))$
$\leq$ $\ T\varepsilon d\ln(eT/d)$
$\in$ $\ O\left(\frac{d}{\sigma}T\varepsilon\log(T+1/\varepsilon)\right)$
@@
Consider an arbitrary $t\in[T]$. If $t\in Q_{T}$, then $\mu_{t}(x_{t},y_{t})=\mu_{t}^{m}(x_{t})$ trivially, so assume $t\not\in Q_{T}$. Let $(x^{\prime},y^{\prime})=\operatorname*{arg\,min}_{(x,y)\in S_{t}:\pi_{t}(x_{t})=y}||x_{t}-x||$. Since $t\not\in Q_{T}$, we must have $||x_{t}-x^{\prime}||\leq\varepsilon^{1/n}$.
We have $y^{\prime}=\pi^{m}(x^{\prime})$ by construction of $S_{t}$ and $\pi_{t}(x_{t})=y^{\prime}$ by construction of $y^{\prime}$. Combining these with the local generalization assumption, we get
$\mu_{t}(x_{t},y_{t})=$ $\ \mu_{t}(x_{t},\pi_{t}(x_{t}))$
$=$ $\ \mu_{t}(x_{t},\pi^{m}(x^{\prime}))$
$\geq$ $\ \mu_{t}^{m}(x_{t})-L||x_{t}-x^{\prime}||$
$\geq$ $\ \mu_{t}^{m}(x_{t})-L\varepsilon^{1/n}$
@@
Under the conditions of Theorem 5.1, Algorithm 1 satisfies
$\operatorname*{\mathbb{E}}\left[R_{T}^{\times}\right]\in$ $\ O\left(\frac{dL}{\sigma\mu_{0}^{m}}T\varepsilon^{1+1/n}\log(T+1/\varepsilon)\right)$
$\operatorname*{\mathbb{E}}\left[R_{T}^{+}\right]\in$ $\ O\left(\frac{dL}{\sigma}T\varepsilon^{1+1/n}\log(T+1/\varepsilon)\right)$
@@
We first claim that $y_{t}=\pi^{m}(x_{t})$ for all $t\not\in M_{T}$. If $t\in Q_{T}$, the claim is immediate. If not, we have $y_{t}=\pi_{t}(x_{t})$ by the definition of the algorithm and $\pi_{t}(x_{t})=\pi^{m}(x_{t})$ by the definition of $t\not\in M_{T}$. Thus $\mu_{t}(x_{t},y_{t})=\mu_{t}^{m}(x_{t})$ for $t\not\in M_{T}$. For $t\in M_{T}$, Lemma B.2 implies that $\mu_{t}^{m}(x_{t})-\mu_{t}(x_{t},y_{t})\leq L\varepsilon^{1/n}$, so
$R_{T}^{+}=$ $\ \sum_{t\in M_{T}}(\mu_{t}^{m}(x_{t})-\mu_{t}(x_{t},y_{t}))$
$\leq$ $\ \sum_{t\in M_{T}}L\varepsilon^{1/n}$
$=$ $\ |M_{T}|L\varepsilon^{1/n}$ (1)
@@
$R_{T}^{\times}\leq\frac{R_{T}^{+}}{\mu_{0}^{m}/2}\leq\frac{2|M_{T}|L\varepsilon^{1/n}}{\mu_{0}^{m}}$ (2)
Taking the expectation and applying Lemma B.1 to Equations 1 and 2 completes the proof. ∎
@@
Let $(K,||\cdot||)$ be a normed vector space and let $\delta>0$. Then a multiset $S\subseteq K$ is a $\delta$-packing of $K$ if for all $a,b\in S$, $||a-b||>\delta$. The $\delta$-packing number of $K$, denoted $\mathcal{M}(K,||\cdot||,\delta)$, is the maximum cardinality of any $\delta$-packing of $K$.
We only consider the Euclidean distance norm, so we just write $M(K,||\cdot||,\delta)=M(K,\delta)$.
Lemma B.4 (Theorem 14.2 in (Wu, 2020)).
@@
If $S\subset\mathbb{R}^{n}$ is compact, then there exists a closed ball with radius at most $\operatorname*{diam}(S)\sqrt{\frac{n}{2(n+1)}}$ containing $S$.
@@
$\operatorname*{\mathbb{E}}[|Q_{T}|]\in O\left(\sqrt{\frac{T}{\varepsilon}}+\frac{d}{\sigma}T\varepsilon\log(T+1/\varepsilon)+\frac{\operatorname*{\mathbb{E}}[\operatorname*{diam}(\boldsymbol{x})^{n}]}{\varepsilon}\right)$
@@
Next, fix a $y\in\mathcal{Y}$ and let $X_{y}=\{x\in\boldsymbol{x}:\pi^{m}(x)=y\}$ be the multiset of observed inputs whose mentor action is $y$. Also let $\hat{X}_{2}=\{x_{t}:t\in\hat{Q}_{2}\}$ be the multiset of inputs associated with time steps in $\hat{Q}_{2}$. Note that $|\hat{X}_{2}|=|\hat{Q}_{2}|$, since $\hat{X}_{2}$ is a multiset. We claim that $\hat{X}_{2}\cap X_{y}$ is a $\varepsilon^{1/n}$-packing of $X_{y}$. Suppose instead that there exists $x,x^{\prime}\in\hat{X}_{2}\cap X_{y}$, with $||x-x^{\prime}||\leq\varepsilon^{1/n}$. WLOG assume $x$ was queried after $x^{\prime}$ and let $t$ be the time step on which $x$ was queried. Then $(x^{\prime},\pi^{m}(x^{\prime}))\in S_{t}$. Also, since $x,x^{\prime}\in\hat{X}_{2}$ we have $\pi_{t}(x_{t})=\pi^{m}(x_{t})=y=\pi^{m}(x^{\prime})$. Therefore
$\min_{(x^{\prime\prime},y^{\prime\prime})\in S_{t}:y^{\prime\prime}=\pi_{t}(x_{t})}||x_{t}-x^{\prime\prime}||\leq||x_{t}-x^{\prime}||\leq\varepsilon^{1/n}$
which contradicts $t\in\hat{Q}$. Thus $\hat{X}_{2}\cap X_{y}$ is a $\varepsilon^{1/n}$-packing of $X_{y}$.
By Lemma B.5, there exists a ball $B_{1}$ of radius $R:=\operatorname*{diam}(\boldsymbol{x})\sqrt{\frac{n}{2(n+1)}}$ which contains $\boldsymbol{x}$. Let $B_{2}$ be the ball with the same center as $B_{1}$ but with radius $\max(R,\varepsilon^{1/n})$. Since $X_{y}\subset\boldsymbol{x}\subset B_{1}\subset B_{2}$ and $\hat{X}_{2}\cap X_{y}$ is a $\varepsilon^{1/n}$-packing of $X_{y}$, $\hat{X}_{2}\cap X_{y}$ is also a $\varepsilon^{1/n}$-packing of $B_{2}$. Also, $B_{2}$ must contain a ball of radius $\varepsilon^{1/n}$, so Lemma B.4 implies that
@@
$=$ $\ \max\left(\operatorname*{diam}(\boldsymbol{x})^{n}\left(\frac{n}{2(n+1)}\right)^{n/2},\>\varepsilon\right)\frac{3^{n}}{\varepsilon}$